OpenStudy (anonymous):

Given two rectangles, one with length = l and width = w, and the second rectangle with length = 2l and width = 4w. How much larger will the perimeter of the second rectangle be than the perimeter of the first rectangle? P2 = P1 + 4w P2 = 2P1 + w P2 = 2P1 + 4w P2 = P1 + w

OpenStudy (anonymous):

p2=2L+4w

OpenStudy (anonymous):

mark o. you mean P2 = 2P1 + 4w

OpenStudy (anonymous):

oh, i didnt noticed its pi,,lol..hold on ..ill fig it out a sec..lol

OpenStudy (anonymous):

yes its p2=2p1+4w ans

OpenStudy (anonymous):

Calculate the volume of a sphere where the radius is 9 meters.

OpenStudy (anonymous):

heres the sol. PERIMETER P=2L+2W L1=L, W1=W...P1=2L+2W L2=2L, W2=4W....P2=2(2L)+2(4W) P2=4L+8W SO P2=2P1+4W P2=2(2L+2W)+4W=4L+4W+4W=4L+8W ANS..

OpenStudy (anonymous):

what is it in cubic meters

OpenStudy (anonymous):

????

OpenStudy (anonymous):

Calculate the volume of a sphere where the radius is 9 meters. VOLUME OF SPHERE V=(4/3)PI*R^3 V=(4/3)PI*(9)^3=972PI ANS..

OpenStudy (anonymous):

V=972PI CU.METER =972*3.14=3053.63 M^3

OpenStudy (anonymous):

did you understand the solution how to get it? any question?